OpenStudy (anonymous):

Distribute $3700 among Joe, Tom and Jack so that joe receives $250 more than Jack and Jack recieves twice hat tom Receives. Indicate the amount that each will receive

OpenStudy (emilylauren):

due to the information given you can come up with 3 equations: Joe: y=x+250 Tom: y=x/2 Jack y=x this is the amount they are given in comparison to one another

OpenStudy (emilylauren):

altogether is amounts to $3,700 so we add the equations together

OpenStudy (emilylauren):

(x+250)+(x/2)+(x)=3,700 then I'm going to multiply everything by two to get rid of the denominator it becomes: 2x+500+x+2x=7,400

OpenStudy (emilylauren):

then combine like terms... 5x+500=7,400 5x=6,900 so x=1380

OpenStudy (emilylauren):

you then take this x amount and plug it into the first three equations for Joe, Tom, and Jack Joe: $1,630 Tom:$690 Jack:$1,380 And that is your final answer